OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x

OpenStudy (anonymous):

The second equation is \[y+4=5x\] Now we get, \[y=5x-4\] The first equation is \[3x+y=1\] Substituting the second equation, we get \[3x+5x-4=1\] That implies, \[3x+5x=1+4\]or\[3x+5x=5\]or\[8x=5\]or\[x=5/8\]Solved.